Tyler and Paul each have a paper route. Tyler has 7 times as many newspapers to deliver as Paul. After Paul took some of Tyler's papers, each boy had 20 papers to deliver. How many newspapers did Paul take?

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52790)

.
Tyler and Paul each have a paper route. Tyler has 7 times as many newspapers to deliver as Paul.
After Paul took some of Tyler's papers, each boy had 20 papers to deliver.
How many newspapers did Paul take?
~~~~~~~~~~~~~~

 x = # of newspapers Paul  had originally;
7x = # of newspapers Tyler had originally.

y = # of newspapers Paul took of Tyler's papers.


Write equations as you read the problem

    7x -  y = 20    (1)

     x +  y = 20    (2)


To solve the system, add the equations.  You will get

    7x + x  = 40

      8x    = 40

       x    = 40/8 = 5.


From equation (2),  y = 20 - x = 20-5 = 15.


ANSWER.  Paul took 15 newspapers.

Solved.



Answer by greenestamps(13200)



When Paul takes some of Tyler's papers, the total number of papers does not change. So, because in the end each boy had 20 papers, the total number of papers was 40.

Initially, Tyler had 7 times as many papers as Paul. So

x = Paul's original number of papers
7x = Tyler's original number of papers

x+7x=8x=40
x=5

Originally, Paul had x=5 papers and Tyler has 7x=35 papers.

Since they each ended up with 20 papers, the number of papers Paul took from Tyler was 15.

ANSWER: 15